OpenStudy (anonymous):

A mathematics journal has accepted 14 articles for publication. However, due to the budgetary restraints only 7 articles can be published this month. How many specific ways can the journal editor assemble 7 of the 14 articles for publication?

OpenStudy (anonymous):

Can someone help me?

OpenStudy (anonymous):

I know the answer, I just don't know how they got it

OpenStudy (notamathgenius):

What did you get?

OpenStudy (anonymous):

I didn't get it, but it's 17,297,280

OpenStudy (anonymous):

Assuming the order doesn't matter: C(14,7) = 14! / (7! * (14 - 7)!) = 14! / (7! * 7!) = 87178291200 / (5040 * 5040) = 87178291200 / 25401600 = 3,432

OpenStudy (anonymous):

I'm redoing a test. And I don't know how to do it, so, I'm asking HOW do I get 17,297,280